Required Information Problem 10-1A Plant asset costs; depreciation methods LO C1, P1 [The following information applies to thRequired 1A Required 1B Required 2 Required 3 Prepare the journal entry to record the purchase. View transaction list JournalRequired 1A Required 1B Required 2 Required 3 Compute the first-year depreciation expense on the building using the straight-Required 1A Required 1B Required 2 Required 3 Compute the first-year depreciation expense on the land improvements assuming aRequired Information Problem 10-1A Plant asset costs; depreciation methods LO C1, P1 [The following Information applies to th
Answer #1

1) Allocation

Allocation of total Cost Appraised value Percent of total appraised value * Total cost of acquisition Apportioned cost
Building 455400 46% * 840000 386400
Land 306900 31% * 840000 260400
Land improvements 69300 7% * 840000 58800
Vehicles 158400 16% * 840000 134400
Total 990000 100% 840000

2) Journal entry

Date account and explanation Debit Credit
Building 386400
Land 260400
Land improvement 58800
Vehicles 134400
Cash 840000
(To record purchase)

3a) Depreciation on building = (386400-31000/15) = 23693

3b) Depreciation on land improvement = 58800*40% = 23520

4) No